OpenStudy (anonymous):

The sides of a rectangle are 5 cm and 10 cm. The sides are then increased by 50% each. The answer is 125 but I don't understand how to work it out. Thanks!

ganeshie8 (ganeshie8):

im getting 112.5 ?

mathslover (mathslover):

What do you need to do here?

Parth (parthkohli):

area

mathslover (mathslover):

Oh.. I see. If the sides are increased by 50% each Then , the new sides will be : 3l/2 * 3b/2 = 9(lb)/4 9/4 * 50 = 9 * 25/2 = 112.5

mathslover (mathslover):

112.5 seems correct to me too.

OpenStudy (anonymous):

50% is half so the sides'd end up being 7.5 & 15 7.5 x 15 = 112.5

Parth (parthkohli):

i still am not sure why the correct answer is 125

Parth (parthkohli):

perhaps because we didn't interpret the question correctly

mathslover (mathslover):

Modern day mathematics.

Miracrown (miracrown):

Trueeeee

Parth (parthkohli):

lol, that too @Zealous

Miracrown (miracrown):

@ParthKohli it takes time to adjust to such things

Parth (parthkohli):

huehue

mathslover (mathslover):

Why do books provide wrong answers?

Parth (parthkohli):

i'm 300% sure this isn't a book and that it is an online class

mathslover (mathslover):

lol... I appreciate your confidence.

Parth (parthkohli):

ok, I'll reduce that down to 200%. i may even be wrong.

mathslover (mathslover):

Naah... reduce that to 199%

Miracrown (miracrown):

you can't rely on anything in these ''modern days''! So every time you do something, make sure you check it at least more than once.

OpenStudy (anonymous):

I got 112.5 too, relieved I'm not alone! I did it like 10x and got the same answer so I got @_@. Thanks so much for the help guys, I'll report to my maths teacher. Yup, this is from a online maths class. 100% agree...."modern day maths".